• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name J. Wexler

    Log entry time 08:00:39 on August25,2009

    Entry number 285834

    keyword=Shift summary - owl 8/25/2009

    Shift summary:

    12AM - Beam still down related to ion chamber issue.
    12:45 - MCC says the ion chamber issue involved the Compton ion chamber being set to trip at 100 instead of 2000. They changed it and gave us beam.
    1:15 - Changing current in the magnets to the PREX tune, and the target to the thin tantalum.
    1:44 - Beam on, 5 uA, no raster on the 45 deg. angled quartz PREX detector.
    2:53 - Seeing only pedestal data in the PREX detector, called for access.
    4:30 - PREX detector changed to the flat configuration, and the 45 deg. configuration showed a marked malfunction (see entry 285797). The trigger was examined, and the left arm was shut off to prevent overheating related to an AC malfunction earlier in the night.
    5:00 - Resume beam, swept dipole to try and find proper tune.
    5:55 - Called for access in order to change trigger on the PREX detector.
    6:30 - Changed trigger T3 from PREX scintillator to PREX quartz detector. Also changed to the thick tantalum target.
    6:50 - Commenced 5uA current runs on the thick Ta target.
    7:28 - Switch to controlled access for target cooldown.

    Runs taken:
    4303-4309 - 45 deg. PREX configuration tests, 5uA beam, thin tantalum.
    4310-4313 - Flat PREX config. cosmic tests during access thin tantalum.
    4314-4315 - Flat PREX config. with 5uA beam, thin tantalum.
    4316-4320 - Flat PREX config. with 5uA beam, thick tantalum.